# SemGauss-SLAM: Dense Semantic Gaussian Splatting SLAM

### v3

@@ 4.1 Experimental Setup
Depth L1[cm] $\downarrow$ ATE RMSE[cm] $\downarrow$ PSNR(dB) $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NICE-SLAM zhu2022nice 2.97 2.51 24.42 0.809 0.233
Vox-Fusion yang2022vox 2.46 1.47 24.41 0.801 0.236
Co-SLAM wang2023co 1.51 1.06 30.24 0.939 0.252
ESLAM eslam 0.95 0.62 29.08 0.929 0.239
SplaTAM keetha2023splatam 0.73 0.41 34.11 0.968 0.102
SNI-SLAM zhu2023sni 0.77 0.46 29.43 0.935 0.235
@@
Figure 3: Qualitative comparison on rendering quality of our method and baseline. We visualize 4 selected scenes of Replica dataset straub2019replica and details are highlighted with red color boxes. Our method achieves photo-realistic rendering quality and higher completion of reconstruction, especially in areas with rich textural information.
4.2 Experimental Results
SLAM and Rendering Quality Results.   As shown in Tab. 1, our method achieves the highest accuracy in all metrics compared with other radiance field-based SLAM and up to 35% relative increase in reconstruction accuracy. This enhancement is attributed to the incorporation of semantic-informed bundle adjustment, which provides multi-view constraint to achieve joint optimization of
Table 2: We compare our proposed SemGauss-SLAM with other existing radiance field-based SLAM methods on ScanNet dataset dai2017scannet for tracking metric RMSE (cm) $\downarrow$.
@@ 4.2 Experimental Results
NICE-SLAM zhu2022nice 12.00 14.00 10.90 13.40 6.20 11.30
Vox-Fusion yang2022vox 68.84 24.18 27.28 23.30 9.41 30.60
Point-SLAM sandstrom2023point 10.24 7.81 22.16 14.77 9.54 12.90
SplaTAM keetha2023splatam 12.83 10.10 12.08 11.10 7.46 10.71
@@
both camera poses and scene representation. Tab. 2 demonstrates our method outperforms baseline methods on real-world dataset ScanNet dai2017scannet .
Novel View Semantic Evaluation Results.   As shown in Tab. 3, our method achieves up to 49% relative increase in semantic novel view synthesis compared with existing dense semantic SLAM. By introducing 3D Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling. This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring semantic representation accuracy from different viewpoints for high-precision novel view semantic synthesis.
Table 3: Quantitative comparison of semantic novel view synthesis performance on Replica straub2019replica for semantic metric mIoU(%). For one scene, we randomly choose 100 new viewpoints that are different from SLAM mapping perspectives for evaluation and mIoU is calculated between splatted semantic labels and ground truth labels. Our method outperforms baseline across all scenes, showing excellent 3D semantic mapping accuracy.
Methods room0 room1 room2 office0 office1 office2 office3 office4 Avg.
SNI-SLAM zhu2023sni 51.20 50.10 54.80 70.21 63.47 58.91 64.31 71.04 60.51
SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 89.57 92.80 92.41 92.10 90.17
Table 4: Quantitative comparison of input views semantic segmentation performance on 4 scenes of Replica straub2019replica for semantic metric mIoU(%), as its results are only reported on such scenes. For one scene, we calculate mIoU between splatted semantic maps and ground truth labels every 4 frames to obtain average mIoU. Our method achieves the highest semantic accuracy across all scenes.
@@
GT NIDS-SLAM haghighi2023neural 82.45 84.08 76.99 85.94 82.37
DNS SLAM li2023dns 88.32 84.90 81.20 84.66 84.77
SNI-SLAM zhu2023sni 88.42 87.43 86.16 87.63 87.41
SGS-SLAM li2024sgs 92.95 92.91 92.10 92.90 92.72
@@
Figure 4: Qualitative comparison on semantic novel view synthesis of our method and baseline semantic SLAM method SNI-SLAM zhu2023sni on 4 scenes of Replica straub2019replica .
@@
Semantic Segmentation Results.   Since haghighi2023neural ; li2023dns ; li2024sgs ; zhu2023sni employs ground truth labels for supervision, for a fair comparison, we also present our results using ground truth labels for supervision, denoted as Our (GT). The results of using semantic segmentation results for supervision are Ours. As shown in Tab. 4, our work outperforms existing radiance field-based semantic SLAM methods. Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization. Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map for high-precision semantic representation.
@@
Visualization.   Fig. 3 shows rendering quality comparison of 4 scenes with interesting regions highlighted with colored boxes. For less-frequent observed areas, such as floor and side of sofa, other methods either result in low-quality reconstruction or leave holes. Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed and complete geometric reconstruction results. Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise and complete geometry reconstruction. As shown in Fig. 4, our method achieves superior novel view semantic segmentation accuracy compared with baseline SNI-SLAM zhu2023sni . It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process. Consequently, the ceiling features are poorly modeled in the semantic scene representation. Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.
4.3 Ablation Study
@@
Figure 5: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica straub2019replica .
@@
We perform ablation study on three scenes of Replica dataset straub2019replica in Tab. 5 to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM. Moreover, ablation of semantic-informed BA component is conducted in Tab. 6.
@@
Feature-level Loss.   Tab. 5 shows that incorporating feature-level loss significantly enhances semantic segmentation performance, while having little effect on tracking and geometric reconstruction. This result occurs because feature-level loss influences only the optimization of semantic features, without affecting the optimization of geometry and pose estimation. As shown in Fig. 5, utilizing feature-level loss can achieve improved boundary segmentation and finer segmentation of small objects. This enhancement occurs because feature loss compels the scene representation to capture high-dimensional and direct information within the feature space, leading to the capability of distinguishing intricate details and subtle variations within the scene.
@@
Semantic-informed BA.   Tab. 5 shows that introducing semantic-informed BA leads to enhancements in tracking, reconstruction, and semantic segmentation. This improvement is due to the joint optimization of camera poses and scene representation, which is informed by multi-view constraints. As shown in Tab. 6, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints. Such result suggests that multi-view semantic constraints provide more comprehensive and accurate information, due to the consistency of semantics across multiple perspectives. Moreover, the absence of semantic constraint can lead to reduced semantic precision, while lacking color and depth constraints results in decreased reconstruction accuracy.
Table 5: Ablation study of our contributions on three scenes of Replica straub2019replica .
Methods room0 room1 office1
RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
w/o feature-level loss 0.26 0.54 83.60 0.42 0.46 86.10 0.17 0.22 80.13
w/o semantic-informed BA 0.35 0.70 90.10 0.51 0.60 91.08 0.21 0.30 87.73
SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
@@
Table 6: Ablation study of semantic-informed BA on three scenes of Replica straub2019replica . (w/o semantic) without adding semantic constraint loss $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding RGB and depth constraint loss $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
Methods room0 room1 office1
RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
w/o semantic 0.35 0.66 90.15 0.50 0.49 91.50 0.20 0.24 87.95
w/o RGB and depth 0.31 0.70 92.01 0.48 0.59 93.90 0.18 0.28 89.60
SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.42 0.46 94.10 0.17 0.22 90.11
5 Conclusion

### v4

@@ IV-A Experimental Setup
[…]
NICE-SLAM [23] 2.97 2.51 24.42 0.809 0.233
Vox-Fusion [24] 2.46 1.47 24.41 0.801 0.236
Co-SLAM [26] 1.51 1.06 30.24 0.939 0.252
ESLAM [25] 0.95 0.62 29.08 0.929 0.239
SplaTAM [12] 0.73 0.41 34.11 0.968 0.102
SNI-SLAM [7] 0.77 0.46 29.43 0.935 0.235
@@
IV-B Experimental Results
SLAM and Rendering Quality Results.   As shown in Tab. I, our method achieves the highest accuracy in all metrics compared with other radiance field-based SLAM and up to 35% relative increase in reconstruction accuracy. This enhancement is attributed to the incorporation of semantic-informed bundle adjustment, which provides multi-view constraint to achieve joint optimization of both camera poses and scene representation. Tab. II demonstrates our method outperforms baseline methods on real-world dataset ScanNet [37]. Note that each scene is tested and averaged with five independent runs to ensure the reliability of results.
TABLE II: Accuracy comparison on ScanNet dataset [37] for tracking metric RMSE (cm) $\downarrow$.
@@ IV-B Experimental Results
NICE-SLAM [23] 12.00 14.00 10.90 13.40 6.20 11.30
Vox-Fusion [24] 68.84 24.18 27.28 23.30 9.41 30.60
Point-SLAM [27] 10.24 7.81 22.16 14.77 9.54 12.90
SplaTAM [12] 12.83 10.10 12.08 11.10 7.46 10.71
@@
Novel View Semantic Evaluation Results.   Tab. III shows that our method achieves up to 49% relative increase in semantic novel view synthesis compared with SNI-SLAM [7], showing excellent 3D semantic mapping accuracy. Note that for one scene, we randomly choose 100 new viewpoints that are different from SLAM mapping perspectives for evaluation and mIoU is calculated between splatted semantic labels and ground truth labels. By introducing Gaussian feature embedding for semantic representation, our method enables continuous semantic modeling. This modeling is crucial for generating semantically coherent scenes, as it reduces the occurrence of sharp transitions that can cause inconsistencies, ensuring accurate semantic representation from novel viewpoints.
TABLE III: Quantitative comparison of semantic novel view synthesis performance on Replica [36] for semantic metric mIoU(%).
@@
SNI-SLAM [7] 51.20 50.10 54.80 70.21 56.58
SemGauss-SLAM (Ours) 89.63 84.72 86.51 93.60 88.62
TABLE IV: Input views semantic segmentation performance on 4 scenes of Replica [36] for semantic metric mIoU(%).
Methods room0 room1 room2 office0 Avg.
GT NIDS-SLAM [9] 82.45 84.08 76.99 85.94 82.37
DNS SLAM [8] 88.32 84.90 81.20 84.66 84.77
SNI-SLAM [7] 88.42 87.43 86.16 87.63 87.41
SGS-SLAM [16] 92.95 92.91 92.10 92.90 92.72
@@
Semantic Segmentation Results.   Since [9, 8, 16, 7] employs ground truth labels for supervision, for a fair comparison, we also present our results using ground truth labels for supervision, denoted as Our (GT). The results of using semantic segmentation results for supervision are Ours. As shown in Tab. IV, our work outperforms existing radiance field-based semantic SLAM methods. Such enhancement attributes to the integration of semantic feature embedding into 3D Gaussian for enriched semantic representation, and semantic feature-level loss for direct guidance of semantic optimization. Moreover, our proposed semantic-informed BA also contributes to high semantic precision, as it leverages multiple co-visible frames to construct a globally consistent semantic map for high-precision semantic representation.
@@
Visualization.   Fig. 3 shows rendering quality comparison of 4 scenes with interesting regions highlighted with colored boxes. For less-frequent observed areas, such as floor and side of sofa, other methods either result in low-quality reconstruction or leave holes. Our method leverages the high-quality rendering capability of 3D Gaussian representation and introduces semantic-informed BA to achieve detailed and complete geometric reconstruction results. Specifically, our method enhances multi-view geometry consistency by BA to ensure that the reconstructed geometry aligns accurately across all observed viewpoints, leading to precise and complete geometry reconstruction. As shown in Fig. 4, our method achieves superior novel view semantic segmentation accuracy compared with baseline SNI-SLAM [7]. It can be observed that SNI-SLAM struggles with segmenting ceilings in novel view synthesis as they are less frequently observed during the mapping process. Consequently, the ceiling features are poorly modeled in the semantic scene representation. Our method introduces semantic-informed BA to construct multi-view constraints, enabling areas that are sparsely observed to effectively utilize the limited information from co-visible frames, thus establishing sufficient constraints for accurate semantic reconstruction.
@@
Figure 5: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica [36].
TABLE V: Ablation study of our contributions on Replica [36].
Methods room0 office1
RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
w/o feature-level loss 0.26 0.54 83.60 0.17 0.22 80.13
w/o semantic-informed BA 0.35 0.70 90.10 0.21 0.30 87.73
SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.17 0.22 90.11
IV-C Ablation Study
@@
We perform ablation study on two scenes of Replica dataset [36] in Tab. V to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM. Moreover, ablation of semantic-informed BA component is conducted in Tab. VI.
@@
Feature-level Loss.   Tab. V shows that incorporating feature-level loss significantly enhances semantic segmentation performance, while having little effect on tracking and geometric reconstruction. This result occurs because feature-level loss influences only the optimization of semantic features, without affecting the optimization of geometry and pose estimation. As shown in Fig. 5, utilizing feature-level loss can achieve improved boundary segmentation and finer segmentation of small objects. This enhancement occurs because feature loss compels the scene representation to capture high-dimensional and direct information within the feature space, leading to the capability of distinguishing intricate details and subtle variations within the scene.
@@
Semantic-informed BA.   Tab. V shows that introducing semantic-informed BA leads to enhancements in tracking, reconstruction, and semantic segmentation. This improvement is due to the joint optimization of camera poses and scene representation, which is informed by multi-view constraints. As shown in Tab. VI, lacking semantic constraint leads to a significant decrease in tracking performance compared with the absence of color and depth constraints. Such result suggests that multi-view semantic constraints provide more comprehensive and accurate information, due to the consistency of semantics across multiple perspectives. Moreover, the absence of semantic constraint can lead to reduced semantic precision, while lacking color and depth constraints results in decreased reconstruction accuracy.
@@
TABLE VI: Ablation study of semantic-informed BA on Replica [36]. (w/o semantic) without adding $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
Methods room0 office1
RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$ RMSE $\downarrow$ Depth L1 $\downarrow$ mIoU $\uparrow$
w/o semantic 0.35 0.66 90.15 0.20 0.24 87.95
w/o RGB and depth 0.31 0.70 92.01 0.18 0.28 89.60
SemGauss-SLAM (Ours) 0.26 0.54 92.81 0.17 0.22 90.11
V Conclusion